Find y'. Simplify all answers.

a. x²y + xy² - y³ + x² = 2

b. sqrt(2xy) + y²-xy²=1

c. sinx + cosy = sinxcosy

d. y = x^x^x

e. y = (sinx)^2x

f. ln(x²+y²) + 2arctan(y/x) = 1

g. (x+y)³ = (x-y)²


1
Expert's answer
2021-11-22T07:33:30-0500

a.


"(x^2y + xy^2- y^3+ x^2 )' = (2)'"

"2xy+x^2y'+y^2+2xyy'-3y^2y'+2x=0"

"y'=\dfrac{2xy+y^2+2x}{3y^2-2xy-x^2}"

b.


"(\sqrt{2xy} + y^2-xy^2)'=(1)'"

"\sqrt{\dfrac{y}{2x}}+\sqrt{\dfrac{x}{2y}}y'+2yy'-y^2-2xyy'=0"

"y'=\dfrac{y^2-\sqrt{\dfrac{y}{2x}}}{\sqrt{\dfrac{x}{2y}}+2y-2xy}"

c.


"(\sin x + \cos y )'= (\sin x\cos y)'"

"\cos x-y'\sin y=\cos x\cos y-y'\sin x \sin y"

"y'=\dfrac{\cos x\cos y-\cos x}{\sin x \sin y-\sin y}"

d.


"y = x^{x^x}"

"\ln y=x^x\ln x"

"\dfrac{\ln y}{\ln x}=x^x"

"\ln(\dfrac{\ln y}{\ln x})=x\ln x"

"(\ln(\ln y)-\ln(\ln x))'=(x\ln x)'"

"\dfrac{1}{y\ln y}y'-\dfrac{1}{x\ln x}=\ln x+1"

"y'=y\ln y(\dfrac{1}{x\ln x}+\ln x+1)"

"y'=x^{x^x}x^x\ln x(\dfrac{1}{x\ln x}+\ln x+1)"

e.


"y = (\sin x)^{2x}"

"(\ln y)'=(2x\ln(\sin x))'"

"\dfrac{y'}{y}=2\ln(\sin x)+2x(\dfrac{\cos x}{\sin x})"

"y'=2(\sin x)^{2x}(\ln(\sin x)+x\cot x)"

f)


"(\ln(x²+y²) + 2\arctan(y/x))' = (1)'"

"\dfrac{2x+2yy'}{x^2+y^2}+\dfrac{2}{1+y^2/x^2}(\dfrac{y'}{x}-\dfrac{y}{x^2})=0"

"x+yy'+xy'-y=0"

"y'=\dfrac{y-x}{y+x}"

g)


"((x+y)^3)' =( (x-y)^2)'"

"3(x+y)^2(1+y')=2(x-y)(1-y')"

"y'=\dfrac{2(x-y)-3(x+y)^2}{3(x+y)^2+2(x-y)}"